# Can 3D point cloud data improve automated body condition score prediction in dairy cattle?

### v1

@@ Introduction
The data used in this study were collected from a commercial dairy farm in central Georgia, USA. Dairy cows were imaged in November 2024 and February 2025 using a depth-sensing image acquisition system equipped with a RealSense D455 camera (Intel Corp., Santa Clara, CA, USA). The camera was installed in a chute system at the exit of the milking parlor, enabling fully automated image acquisition as cows returned to the pen after milking. Each cow was independently scored by two trained evaluators using a five-point BCS scale with 0.25 point increments (Ferguson et al., 1994). Only cows with exact agreement between the two scorers were retained for modeling. In total, 1,020 dairy cows, comprising 10,228 depth CSV files and corresponding RGB images, were included in the study. These cows ranged in age from 3 to 10 years and had BCS values from 2.0 to 4.25. All animal procedures conducted for data collection were approved by the University of Florida Institutional Animal Care and Use Committee (IACUC protocol number 202200000534).
@@
A keypoint detection model was developed to identify anatomical landmarks for experimental settings 3 and 4, which were manually annotated on RGB images using LabelMe. These include, left and right short rib points, left and right hook points, and left and right pin points. To predict these seven anatomical keypoints across all 1,020 cows, a keypoint detection model based on YOLOv11n-pose was trained using a randomly selected subset of 877 cows, with one image per cow. These images were randomly partitioned into 70% for training, 15% for validation, and 15% for testing to develop and evaluate the model. The YOLOv11n-pose model was trained with geometric and photometric augmentations implemented using the Albumentations library (Buslaev et al., 2020). These augmentations included horizontal flipping (probability 0.5), affine transformations (probability 0.7; translation up to 2% of image size, isotropic scaling within $\pm 5\%$, and rotation within $\pm 10^{\circ}$), and appearance perturbations (probability of 0.2 each for brightness/contrast adjustment and additive Gaussian noise). Keypoint detection accuracy was evaluated using the percentage of correct keypoints (PCK) and root mean square error (RMSE). PCK measures the proportion of predicted keypoints whose Euclidean distance from the ground-truth location falls within a specified pixel threshold (k), providing a tolerance-based assessment of accuracy. RMSE quantifies the average localization error in pixels across all keypoints. Together, these metrics characterize both tolerance-dependent accuracy and absolute localization precision of anatomical landmark predictions. Once satisfactory keypoint detection performance was achieved, the trained model was used to automatically detect keypoints for all 1,020 cows across 10,228 RGB images.
@@
A total of six predicted anatomical keypoints from the trained pose model were refined to improve spatial accuracy in three-dimensional space (Figure 1). The hook and pin keypoints were adjusted by searching for the highest local point within a defined neighborhood around the initial predictions (30 pixels for hooks and 10 pixels for pins). Based on these refined landmarks, additional three anatomical keypoints were computed: spike A (the midpoint between the left and right short ribs), spike B (the midpoint between the adjusted left and right hooks), and spike C (the midpoint between the adjusted left and right pins) (Figure S1), resulting in a total of nine anatomical keypoints. Three classes of handcrafted features were then extracted from both depth images and point clouds: maximum distance, area, and volume.
@@
The volume was partitioned into four regions (V1, V2, V3, and V4) (Figure 2). Using V1 as an illustration for depth images, V1 represents the total volume of the cow’s body surface located above the anatomical plane defined by Spike A, Spike B, and the right hook point (Figure 2). This volume was calculated by integrating the vertical distance between the body surface and the reference plane over the corresponding triangular region. Larger values indicate greater body mass distribution in this area.
For point cloud–based handcrafted features, depth rasters were first converted into calibrated three-dimensional point clouds. The same anatomical keypoints and 10 lines derived from the depth images were mapped into 3D space. Maximum distance, area, and volume features were then defined (Figure 3). This parallel feature extraction strategy ensured consistency between depth image and point cloud representations while capturing geometric characteristics in true three-dimensional space.
@@
Hyperparameters were optimized using a Bayesian optimization search implemented in Optuna. The search space included the learning rate, weight decay, dropout rate, fully connected layer size, batch size, image resolution, number of training epochs, and early stopping patience (Table 1). For each trial, the model was trained on the training set and evaluated on the validation set, with validation accuracy used as the optimization objective, while validation loss controlled early stopping and learning rate scheduler updates. Random seeds were fixed across Python, NumPy, and PyTorch to ensure deterministic and reproducible results. After hyperparameter optimization, the best performing configuration was retrained from scratch using the combined training and validation sets and the final model performance was evaluated on the test set. This workflow was applied identically to ResNet-18 and ConvNeXt-Tiny, enabling a consistent and fair comparison. Both ResNet-18 and ConvNeXt-Tiny were implemented using the Torchvision Python package (Paszke et al., 2019).
@@
Model performance was evaluated using cow-level random subsampling cross-validation repeated five times to assess generalization to unseen animals and prevent data leakage (Wang et al., 2025). Final performance was reported as the mean and standard deviation of cow-level metrics across repeats. In each evaluation run, all images belonging to a given cow were assigned exclusively to either the training (70%), validation (15%), or testing (15%) partition, ensuring that no cow appeared in more than one subset. This design eliminates identity leakage, which can otherwise artificially inflate performance when multiple images from the same animal are included across partitions. Within each training partition, a validation subset was created at the cow level and used for early stopping, learning rate scheduling, and hyperparameter tuning. Class distributions were monitored during partitioning, and cows were allocated to subsets to approximate stratification by BCS category while maintaining strict cow-level separation. All model hyperparameters were finalized prior to testing and were not adjusted using the test data.

### v2

@@ Introduction
The data used in this study were collected from a commercial dairy farm in central Georgia, USA. Dairy cows were imaged in November 2024 and February 2025 using a depth-sensing image acquisition system equipped with a RealSense D455 camera (Intel Corp., Santa Clara, CA, USA). The camera was installed 2515 mm above the ground in a chute system at the exit of the milking parlor, enabling fully automated image acquisition as cows returned to the pen after milking. Each cow was independently scored by two trained evaluators using a five-point BCS scale with 0.25 point increments (Ferguson et al., 1994). Only cows with exact agreement between the two scorers were retained for modeling. In total, 1,020 dairy cows, comprising 10,228 depth CSV files and corresponding RGB images, were included in the study. These cows ranged in age from 3 to 10 years and had BCS values from 2.0 to 4.25. All animal procedures conducted for data collection were approved by the University of Florida Institutional Animal Care and Use Committee (IACUC protocol number 202200000534).
@@
A keypoint detection model was developed to identify anatomical landmarks for experimental settings 3 and 4, which were manually annotated on RGB images using LabelMe. These include, left and right short rib points, left and right hook points, and left and right pin points. To predict these six anatomical keypoints across all 1,020 cows, a keypoint detection model based on YOLOv11s-pose was trained using a randomly selected subset of 877 cows, with one image per cow. These images were randomly partitioned into 70% for training, 20% for validation, and 10% for testing to develop and evaluate the model. The YOLOv11s-pose model was trained with geometric and photometric augmentations implemented using the Albumentations library (Buslaev et al., 2020). These augmentations included horizontal flipping (probability 0.5), affine transformations (probability 0.7; translation up to 2% of image size, isotropic scaling within $\pm 5\%$, and rotation within $\pm 10^{\circ}$), and appearance perturbations (probability of 0.2 each for brightness/contrast adjustment and additive Gaussian noise). Keypoint detection accuracy was evaluated using the percentage of correct keypoints (PCK) and root mean square error (RMSE). PCK measures the proportion of predicted keypoints whose Euclidean distance from the ground-truth location falls within a specified pixel threshold (k), providing a tolerance-based assessment of accuracy. RMSE quantifies the average localization error in pixels across all keypoints. Together, these metrics characterize both tolerance-dependent accuracy and absolute localization precision of anatomical landmark predictions. Once satisfactory keypoint detection performance was achieved, the trained model was used to automatically detect keypoints for all 1,020 cows across 10,228 RGB images.
@@
A total of six predicted anatomical keypoints from the trained YOLOv11s-pose model were refined to improve spatial accuracy in three-dimensional space (Figure 1). The hook and pin keypoints were adjusted by searching for the highest local point within a defined neighborhood around the initial predictions (30 pixels for hooks and 10 pixels for pins). Based on these refined landmarks, additional three anatomical keypoints were computed: spike A (the midpoint between the left and right short ribs), spike B (the midpoint between the adjusted left and right hooks), and spike C (the midpoint between the adjusted left and right pins) (Figure S1), resulting in a total of nine anatomical keypoints. Three classes of handcrafted features were then extracted from both depth images and point clouds: maximum distance, area, and volume.
@@
The volume was partitioned into four regions (V1, V2, V3, and V4) (Figure 2). Using V1 as an illustration for depth images, V1 represents the total volume enclosed between the cow’s body surface and the anatomical plane defined by spike A, spike B, and the right hook point (Figure 2). This volume was calculated by summing the vertical difference between the heights of the anatomical plane and the cow’s body surface over all points within the triangular region, with both heights measured relative to the ground (Zhao et al., 2023). To estimate the height of all points within the triangular anatomical plane, the coordinates ($x$, $y$) and corresponding height ($z$) of spike A, spike B, and the right hook point were used to determine the plane equation $z=ax+by+c$ and obtain $a$, $b$, and $c$ using the NumPy package (Harris et al., 2020). These estimates were then used to compute the height of a given point ($x$, $y$) on the anatomical plane for volume calculation. Larger volume values indicate lower body mass distribution in this area.
For point cloud–based handcrafted features, depth rasters were first converted into calibrated three-dimensional point clouds. The same anatomical keypoints and 10 lines derived from the depth images were mapped into 3D space. Maximum distance, area, and volume (Zhao et al., 2023) features were then defined (Figure 3). This parallel feature extraction strategy ensured consistency between depth image and point cloud representations while capturing geometric characteristics in true three-dimensional space.
@@
Hyperparameters were optimized using a Bayesian optimization search implemented in Optuna (Akiba et al., 2019). The search space included the learning rate, weight decay, dropout rate, fully connected layer size, batch size, image resolution, number of training epochs, and early stopping patience (Table 1). For each trial, the model was trained on the training set and evaluated on the validation set, with validation accuracy used as the optimization objective, while validation loss controlled early stopping and learning rate scheduler updates. Random seeds were fixed across Python, NumPy, and PyTorch to ensure deterministic and reproducible results. After hyperparameter optimization, the best performing configuration was retrained from scratch using the combined training and validation sets and the final model performance was evaluated on the test set. This workflow was applied identically to ResNet-18 and ConvNeXt-Tiny, enabling a consistent and fair comparison. Both ResNet-18 and ConvNeXt-Tiny were implemented using the Torchvision Python package (Paszke et al., 2019).
@@
Model performance was evaluated using cow-level random subsampling cross-validation repeated five times to assess generalization to unseen animals and prevent data leakage (Wang et al., 2025). Final performance was reported as the mean and standard error of cow-level metrics across repeats. In each evaluation run, all images belonging to a given cow were assigned exclusively to either the training (70%), validation (15%), or testing (15%) partition, ensuring that no cow appeared in more than one subset. This design eliminates identity leakage, which can otherwise artificially inflate performance when multiple images from the same animal are included across partitions. Within each training partition, a validation subset was created at the cow level and used for early stopping, learning rate scheduling, and hyperparameter tuning. Class distributions were monitored during partitioning, and cows were allocated to subsets to approximate stratification by BCS category while maintaining strict cow-level separation. All model hyperparameters were finalized prior to testing and were not adjusted using the test data.
@@ Conclusion
Supplemental Material
Figure S1: Keypoint detection and derived spike landmarks. The figure shows nine anatomical keypoints identified by the keypoint detection model. Derived spike landmarks were computed from these keypoints and used as reference points for downstream feature extraction and quantitative analysis.
Figure S2: Distribution of body condition score data. Observations collected in 2024 and 2025 are shown in different colors in the histogram.
Figure S3: F1–confidence curve for the YOLO segmentation model on the test set. The F1 score (y-axis) was evaluated across different confidence thresholds (x-axis). The thin line represents the performance for the cow instance, while the thick line indicates the smoothed average performance across confidence thresholds. The maximum F1 score (1.00) is achieved at a confidence threshold of 0.925, which was selected as the optimal operating point for subsequent evaluations.
Table S1: A list of handcrafted features derived from depth images and point cloud data that were used in prediction models in experimental setting 4. MaxDist refers to the maximum distance.
Feature type Feature name Description of handcrafted feature
Maximum distance MaxDist (L1) Maximum perpendicular distance between the cow body surface curve and line L1 connecting spike A and the right hook point
MaxDist (L2) Maximum perpendicular distance between the cow body surface curve and L2 connecting Spike B and the right hook point
MaxDist (L3) Maximum perpendicular distance between the cow body surface curve and line L3 connecting Spike C and the right hook point
MaxDist (L4) Maximum perpendicular distance between the cow body surface curve and line L4 connecting right pin and the right hook point
MaxDist (L5) Maximum perpendicular distance between the cow body surface curve and line L5 connecting spike B and right pin point
MaxDist (L6) Maximum perpendicular distance between the cow body surface curve and line L6 connecting spike B and left pin point
MaxDist (L7) Maximum perpendicular distance between the cow body surface curve and line L7 connecting left hook and left pin point
MaxDist (L8) Maximum perpendicular distance between the cow body surface curve and line L8 connecting left hook and spike C point
MaxDist (L9) Maximum perpendicular distance between the cow body surface curve and line L9 connecting left hook and spike B point
MaxDist (L10) Maximum perpendicular distance between the cow body surface curve and line L10 connecting left hook and spike A point
Area Area (L1) Area enclosed between the cow body surface curve and line L1 connecting spike A and the right hook point
Area (L2) Area enclosed between the cow body surface curve and line L2 connecting spike B and the right hook point
Area (L3) Area enclosed between the cow body surface curve and line L3 connecting spike C and the right hook point
Area (L4) Area enclosed between the cow body surface curve and line L4 connecting right pin and the right hook point
Area (L5) Area enclosed between the cow body surface curve and line L5 connecting spike B and right pin point
Area (L6) Area enclosed between the cow body surface curve and line L6 connecting spike B and left pin point
Area (L7) Area enclosed between the cow body surface curve and line L7 connecting left hook and left pin point
Area (L8) Area enclosed between the cow body surface curve and line L8 connecting left hook and spike C point
Area (L9) Area enclosed between the cow body surface curve and line L9 connecting left hook and spike B point
Area (L10) Area enclosed between the cow body surface curve and line L10 connecting left hook and spike A point
Volume Volume (V1) Volume enclosed between the cow body surface and the triangular plane defined by spike A, spike B, and the right hook point
Volume (V2) Volume enclosed between the cow body surface and the triangular plane defined by spike A, spike B, and the left hook point
Volume (V3) Volume enclosed between the cow body surface and the triangular plane defined by spike B, spike C, and the right hook point
Volume (V4) Volume enclosed between the cow body surface and the triangular plane defined by spike B, spike C, and the left hook point